michael saved 50 php in january. suppose he saved twice the amount he saved the previous month. how much would he save at the end of one year ? Use the geometric series technique
Answer by jorel555(1290)

50+100+200+400+...=50((2^n)-1)where n is the number of periods. Since n, in this case, is 12, we have:
50x((2^12)-1)=50x4095=204,800 php saved in one year!!!!!!!!!!!